# Application of Laplace filters to the analysis of lattice time correlators

### v1

@@ I Introduction
Lattice quantum chromodynamics (QCD), or more generally, lattice field theory, allows one to formulate strongly coupled field theories nonperturbatively. Such a formulation is then often used as the basis for numerical ab initio simulations, which provide systematically improvable predictions in particle physics and related fields. Compelling examples of results obtained using lattice QCD can be found in the Flavour Lattice Averaging Group (FLAG) review [1] and the Muon g2g-2 Theory Initiative White Paper [2]. In lattice QCD, a discretized version of the QCD path integral in Euclidean space-time is numerically estimated using Markov chain Monte Carlo (MCMC) techniques. The quantities typically computed are correlation functions, i.e., expectation values of products of local quantum operators. They are generally expressed as functions of the momenta and Euclidean times of the operators involved (time-momentum representation). Assuming all time separations between operators are fixed except one, and assuming a theory with a mass gap $M>0$, correlation functions are linear combinations of exponentials of the form $e^{-Et}$, where $t$ is the variable Euclidean time separation, and $E\geq M$ is the energy of a specific quantum state propagating in the system. The coefficients (or amplitudes) of such combinations are related to quantum matrix elements of the operators involved. The fact that lattice correlation functions are finite linear combination of such exponentials is a consequence of the finite number of degrees of freedom in a lattice simulation. The energies and matrix elements are physical quantities typically required for predicting a given observable.
A critical step in most lattice analyses is therefore to model statistical data as linear combinations of exponentials with real exponents. This is generally done through a goodness-of-fit optimization of such a model to the data and is notoriously challenging to achieve, in part due to the two issues below:
@@
The issues above, as well as potential mitigation strategies, have been discussed at great length in the lattice QCD literature, and we review in this introduction some of the most recent works on this topic. One approach to addressing the first issue above is to use the better-behaved Ledoit-Wolf shrinkage estimators [3] for the covariance matrix. Such an estimator replaces the sample covariance matrix $V$ with a regulated version of the form $V^{*}=(1-\lambda)V+\lambda V_{\mathrm{diag}}$, where $V_{\mathrm{diag}}$ is the diagonal part of $V$. It can be proven [3] that the regulator $\lambda$ can be chosen such that, asymptotically, $V^{*}$ is a better-conditioned unbiased estimator of the covariance matrix. This approach is commonly used in lattice QCD; see, for example, [4] as one of the earliest works to use it. A key issue with this approach is that although $V^{*}$ can be shown to be an optimal regularization asymptotically, it may be difficult to estimate the deviation from this behavior at finite statistics. Without studying the asymptotic properties, Ledoit-Wolf shrinkage becomes essentially an ad hoc regularization procedure, and it is difficult to assess its effect on estimating $p$-values and resulting hypothesis tests.
@@
Another approach proposed in [5] suggests using an alternative statistic to the standard correlated $\chi^{2}$, where the covariance matrix is replaced by a better-behaved one, typically the diagonal part of the covariance matrix. Other common regularization strategies such as cuts in the singular value decomposition (SVD) of the correlation matrix can also be included in this formalism. In this work, the authors derive the appropriate $p$-value for this modified $\chi^{2}$, allowing them to perform hypothesis tests based on a more reliable estimator. While this addresses some shortcomings of more ad hoc regularization strategies, the fit parameters obtained by minimizing the modified $\chi^{2}$ are, strictly speaking, not the maximum-likelihood parameters given the correlated distribution of the data.
@@
Finally, regarding the determination of a single-state contribution (issue 2 above), a number of techniques are commonly employed. First, excited-state contamination can be suppressed at the level of the measurement itself by choosing operators with weaker coupling to the unwanted states. This can be achieved, for example, by smearing operators using Jacobi iterations [6] or more general forms targeting specific physics (see, for example, the recent work [7]). Methods based on operator design may not be confined to purely statistical data analysis and may require re-measuring expensive correlation functions. Restricting ourselves to methods usable at the data-analysis stage, one can attempt to use the time-translation properties of correlation functions to reconstruct the spectrum. This is historically the aim of Prony’s method [8], which was adapted to lattice QCD analysis as a generalized eigenvalue problem (GEVP) in [9, 10]. More recently, Wagman proposed a new method [11] using time translations of the correlation function to infer the spectrum of the transfer matrix. This was proven later [12] to be equivalent to the Prony GEVP, and this equivalence is further discussed in [13]. Although methods involving solving for a matrix spectrum have been used consistently in lattice QCD [14], they can feature challenging eigenvalue-identification issues due to the noisy nature of the data. Although there are ways to regularize this problem, including in the previously cited references, one cannot entirely exclude the possibility that systematic biases remain.
@@
In this paper, we propose a new method to solve the two issues above, with an emphasis on simplicity and minimal deviations from standard statistical frameworks. The key concept is to use regulated Laplace filters on correlation functions, which are a discretized version of the second-order differential operator $\partial^{2}_{t}+\lambda^{2}$, where $\lambda$ is an arbitrary real, nonzero regulator that guarantees the invertibility of the filter. Laplace filters are high-pass filters, allowing us to decorrelate data and improve the conditioning of the correlation matrix. Additionally, they annihilate $e^{\pm\lambda t}$ at nonzero times, allowing us to modify and probe the spectrum of a correlation function by varying the regulator. In section˜II, we give general definitions that will be used in the rest of the paper. In section˜III, we introduce representative lattice QCD data that will be used for numerical tests of our methods. In section˜IV, we discuss how Laplace filters can be used to significantly improve the conditioning of covariance matrices. We then demonstrate in section˜V how statistical constraints on the spectrum of a correlation function can be designed using Laplace filters. Finally, we present in section˜VI a number of applications to realistic lattice QCD data-analysis examples.
@@ II General problem and definitions
We consider space-time to be a finite discrete hyper-cubic [^1] This is the most straightforward and common choice in lattice field theory, however every method in this paper can be applied to anisotropic lattices.. We consider an arbitrary quantum expectation value lattice with spacing $a$ and Euclidean metric
@@
By construction, $\mathrm{bar}{\eta}[t]=0$. An unbiased estimator of the time covariance matrix of $C[t]$ is given by
@@
and the associated correlation matrix is defined as
@@
$\Gamma=(D\otimes D)\odot V\,$, (9)
where $\odot$ is the element-wise Hadamard product. Therefore, knowing $(D\otimes D)$ from the data, inverting $V$ can be reduced to inverting $\Gamma$, whose coefficients are bounded in the [1,1][-1,1] interval. We now discuss a metric that can be used to quantify the conditioning of $\Gamma$.
@@ II.1 Correlation dynamic range (CDR)
$\kappa(M)=\frac{\max_{j}(\lambda_{j})}{\min_{j}(\lambda_{j})}\,$, (10)
where the $\lambda_{j}$, $0\leq j<N_{t}$, are the eigenvalues of $M$. The CDR of $M$ is then defined by the value of $\kappa(M)$ in decibels ($\mathrm{d}\mathrm{B}$), namely
@@
The CDR gives the number of orders of magnitude between the smallest and largest eigenvalues of $M$, multiplied by $10$. A general rule of thumb in interpreting the CDR is as follows: a CDR of $x\leavevmode\nobreak\ \mathrm{d}\mathrm{B}$ implies that the precision of the data will be degraded by $x/10$ significant digits when multiplied by $M$.
@@ II.2 General data transformations
We now present the particular type of linear transformation considered in this paper.
@@ III Representative sample data
Regarding the M0M-mes-ll, F1M-mes-lh, and F1M-mes-hh correlation functions, we consider standard operators that induce the quantum numbers of pseudoscalar mesons of the form
@@ IV Improvement of time correlations
In practice, one rarely needs to fit the entire time range of a lattice correlator. Instead, fits generally focus on a limited range where the correlation function $C$ is expected to exhibit asymptotic behavior dictated by its spectral representation. It is therefore useful to define a restriction operator $R_{S}$, where $S$ is a subset of $\{0,\dots,N_{t}-1\}$, that projects $C$ onto the components indexed by $S$. If $S$ is an interval $\{t_{i},\dots,t_{f}-1\}$, we define $R_{t_{i},t_{f}}=R_{S}$. It is clear that $R_{S}$ is a linear operator. We additionally define the following shorthand notation
@@ IV.3 Summary of correlation improvements
In table˜2, we summarize the CDRs obtained by filtering and downsampling the M0M-mes-ll and F1M-mes-lh correlators, for both the full time range and a restricted, representative fit range. In all cases, a combination of downsampling and optimized filtering improves the condition number of the correlation matrix by at least four orders of magnitude. This is an encouraging result, and we emphasize that this improvement is achieved solely through a linear, invertible transformation of the data and does not introduce any additional systematic error in subsequent fits of the correlators.
@@ V Spectral effect of Laplace filters
Figure 5: Representation of the pion correlator M0M-mes-$l$ 2-state test $p$-value $p_{\Lambda}$ defined in eq.˜54, with $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$. The four different plots correspond to various values of the initial time $t_{i}$. In all cases, the final time is fixed to $t_{f}=48$. Values below $1\%$ are not represented (2-state hypothesis rejected at $99\%$ confidence level), and values above $1\%$ are within the solid black contour. Values between $1\%$ and $5\%$ are represented in red, and values above $5\%$ are within the dashed contour, and represented by a gradient between orange and green.
Figure 6: Representation of the heavy-light correlator F1M-mes-lh 2-state test $p$-value $p_{\Lambda}$ defined in eq.˜54, with $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$. The four different plots correspond to various values of the initial time $t_{i}$. In all cases, the final time is fixed to $t_{f}=36$. Plotting conventions are identical to fig.˜5.
@@ V.1.2 Discrete and periodic time
eq.˜42 is the discrete and periodic equivalent of the Källén-Lehmann spectral representation eq.˜34. For $t\neq 0\ (\mathrm{mod}\ N_{t})$, we have
@@
An important conclusion from eq.˜45 is that the Laplace filter can entirely remove the contribution of a given state from a correlator if the regulator is tuned to the energy $\tilde{E}$ associated with this state. On the one hand, this means that when choosing the regulator, one must be careful to avoid the energies of states of interest. On the other hand, it also implies that the filter can be exploited to remove unwanted states, typically excited-state contamination from hadronic interpolating operators.
@@ V.2 Applications to statistical data analysis
Throughout this section, we assume that the correlator under study is given by eq.˜42, with $N_{s}$ states of nonzero amplitude and distinct energies indexed by $n\in\{0,\dots,N_{s}-1\}$. For $t\notin\{-1,0,1\}\ (\mathrm{mod}\ N_{t})$, a single application of the Laplace filter yields eq.˜45. For a vector of $r$ regulators $\Lambda=(\lambda_{0},\dots,\lambda_{r-1})$, we define
@@
i.e., $\smash{C_{\Lambda}}$ is the correlator transformed by $r$ successive Laplace filters. eq.˜45 can be generalized to multiple filters. For $t\notin\{-r,\dots,r\}\ (\mathrm{mod}\ N_{t})$, one obtains
@@
For the remainder of this discussion, we assume that $t\notin\{-r,\dots,r\}\ (\mathrm{mod}\ N_{t})$. In order for the formula above to be valid for at least one time slice, we require $r<N_{t}/2$. Then, eq.˜48 offers an opportunity to determine the energy spectrum $\smash{\tilde{E}_{n}}$ by finding a sequence of Laplace filters that annihilate $C[t]$, i.e., that solve the equation
@@
for all $t$ in a given range satisfying the previous constraints, and containing at least $r$ elements. If the number of filters equals the number of states, i.e., $r=N_{s}$, then eq.˜49 is satisfied if and only if, for all $j\in\{0,\dots,N_{s}-1\}$,
@@
In this case, the energies $E_{n}$ are uniquely determined [^3] This is a consequence of the fact that functions of the form $t\mapsto e^{-Et}$ are linearly independent for distinct values of $E$. by eq.˜49. If $r>N_{s}$, then necessarily the spectrum is a subset of the solution set of eq.˜49, i.e., $\smash{\{\tilde{E}_{0},\dots,\tilde{E}_{N_{s}-1}\}\subset\{\lambda_{0},\dots,\lambda_{r-1}\}}$. If $r<N_{s}$, then eq.˜49 has no solution.
@@
    The number of states $N_{s}$ is given by the minimal $r$ for which eq.˜49 admits a solution;
@@
    The spectrum $\tilde{E}_{j}$ is given by the unique solution of eq.˜49 for $r=N_{s}$.
The procedure above can serve as the basis for new methods to determine the energy spectrum of lattice correlators. Although equation eq.˜49 determines the spectrum uniquely, in practice with noisy simulation data it can only be solved up to statistical noise, introducing ambiguities in the above procedure. Therefore, in what follows, we discuss potential hypothesis-testing methods to constrain the spectrum of the correlator.

### v2

@@ I Introduction
Lattice quantum chromodynamics (QCD), or more generally, lattice field theory, allows one to formulate strongly coupled field theories nonperturbatively. Such a formulation is then often used as the basis for numerical ab initio simulations, which provide systematically improvable predictions in particle physics and related fields. Compelling examples of results obtained using lattice QCD can be found in the Flavour Lattice Averaging Group (FLAG) review [1] and the Muon g2g-2 Theory Initiative White Paper [2]. In lattice QCD, a discretized version of the QCD path integral in Euclidean space-time is numerically estimated using Markov chain Monte Carlo (MCMC) techniques. The quantities typically computed are correlation functions, i.e., expectation values of products of local quantum operators. They are generally expressed as functions of the momenta and Euclidean times of the operators involved (time-momentum representation). Assuming all time separations between operators are fixed except one, and assuming a theory with a mass gap $M>0$, correlation functions are linear combinations of exponentials of the form $e^{-Et}$, where $t$ is the variable Euclidean time separation, and $E\geq M$ is the energy of a specific quantum state propagating in the system. The coefficients (or amplitudes) of such combinations are related to quantum matrix elements of the operators involved. The energies and matrix elements are physical quantities typically required for predicting a given observable.
A critical step in most lattice analyses is therefore to fit statistical data for correlation functions to a truncated spectral decomposition as described above. This is generally done through a goodness-of-fit optimization of such a model to the data. Although this is in appearance a simple problem, precisely assessing systematic uncertainties on the resulting fit parameters can be challenging, in part due to the two issues below:
@@
The issues above, as well as potential mitigation strategies, have been discussed at length in the lattice QCD literature over the past 30 years. Some original work on the topic can be found in Refs. [3, 4], and here we review some of the most recent works on this topic. One approach to addressing the first issue above is to use the better-behaved Ledoit-Wolf shrinkage estimators [5] for the covariance matrix. Such an estimator replaces the sample covariance matrix $V$ with a regulated version of the form $V^{*}=(1-\lambda)V+\lambda V_{\mathrm{diag}}$, where $V_{\mathrm{diag}}$ is the diagonal part of $V$. It can be proven [5] that the regulator $\lambda$ can be chosen such that, asymptotically, $V^{*}$ is a better-conditioned unbiased estimator of the covariance matrix. This approach is commonly used in lattice QCD; see, for example, [6] as one of the earliest works to use it. A key issue with this approach is that although $V^{*}$ can be shown to be an optimal regularization asymptotically, it may be difficult to estimate the deviation from this behavior at finite statistics. The shrinkage described above was then improved using a non-linear optimization in Ref. [7, 8]. However, without studying the asymptotic properties, Ledoit-Wolf shrinkage becomes essentially an ad hoc regularization procedure, and it is difficult to assess its effect on estimating $p$-values and resulting hypothesis tests.
@@
Another approach proposed in [9] suggests using an alternative statistic to the standard correlated $\chi^{2}$, where the covariance matrix is replaced by a better-behaved one, typically the diagonal part of the covariance matrix. This formalism covers other common regularization strategies such as cuts in the singular value decomposition of the correlation matrix (cf. for example Ref. [10]). In this work, the authors derive the appropriate $p$-value for this modified $\chi^{2}$, allowing them to perform hypothesis tests based on a more reliable estimator. While this addresses some shortcomings of more ad hoc regularization strategies, the fit parameters obtained by minimizing the modified $\chi^{2}$ are, strictly speaking, not the maximum-likelihood parameters given the correlated distribution of the data.
@@
Finally, regarding the determination of a single-state contribution (issue 2 above), a number of techniques are commonly employed. First, excited-state contamination can be suppressed at the level of the measurement itself by choosing operators with weaker coupling to the unwanted states. This can be achieved, for example, by smearing operators using Jacobi iterations [11] or more general forms targeting specific physics (see, for example, the recent work [12]). Methods based on operator design may not be confined to purely statistical data analysis and may require re-measuring expensive correlation functions.
Restricting ourselves to methods usable at the data-analysis stage, one can attempt to use the time-translation properties of correlation functions to reconstruct the spectrum. This is historically the aim of Prony’s method [13], which was adapted to lattice QCD analysis as a generalized eigenvalue problem (GEVP) in [14, 15]. More recently, Wagman proposed a new method [16] using time translations of the correlation function to infer the spectrum of the transfer matrix via the Lanczos algorithm. This was proven later [17] to be equivalent to the Prony GEVP, and this equivalence is further discussed in [18]. Although methods involving solving for a matrix spectrum have been used consistently in lattice QCD [19], they can feature challenging eigenvalue-identification issues due to the noisy nature of the data. There are ways to regularize this problem, including in the previously cited references, or additionally in the recent work [20].
@@
In this paper, we propose a new method to solve the two issues above, with an emphasis on simplicity and minimal deviations from standard statistical frameworks. We stress that this method is complementary to the literature reviewed in this section, in the sense that in cases where it does not by itself overcome the mentioned issues, it can be applied in conjunction with other ways to regularise the resulting covariance matrix. The key concept is to use regulated Laplace filters on correlation functions, which are a discretized version of the second-order differential operator $\partial^{2}_{t}+\lambda^{2}$, where $\lambda$ is an arbitrary real, nonzero regulator that guarantees the invertibility of the filter. Laplace filters are high-pass filters, allowing us to decorrelate data and improve the conditioning of the correlation matrix. Additionally, they annihilate $e^{\pm\lambda t}$ at nonzero times, allowing us to modify and probe the spectrum of a correlation function by varying the regulator. In section˜II, we give general definitions that will be used in the rest of the paper. In section˜III, we introduce representative lattice QCD data that will be used for numerical tests of our methods. In section˜IV, we discuss how Laplace filters can be used to significantly improve the conditioning of covariance matrices. We then demonstrate in section˜V how statistical constraints on the spectrum of a correlation function can be designed using Laplace filters. Finally, we present in section˜VI a number of applications to realistic lattice QCD data-analysis examples.
@@ II General problem and definitions
We consider space-time to be a finite discrete hyper-cubic lattice. [^1] This is the most straightforward and common choice in lattice field theory, however every method in this paper can be applied to anisotropic lattices. We consider an arbitrary quantum expectation value lattice with spacing $a$ and Euclidean metric
@@
By construction, $\mathrm{bar}{\eta}[t]=0$. An unbiased estimator of the true time covariance matrix $\Sigma[t_{1},t_{2}]$ of $C[t]$ is given by the sample covariance matrix
@@
and the sample correlation matrix is defined as
@@
$\Gamma=\mathrm{diag}(D)\,V\,\mathrm{diag}(D)\,$, (9)
where $\mathrm{diag}(D)$ is the diagonal matrix with coefficients given by the elements of $D$. Therefore, knowing $(D\otimes D)$ from the data, inverting $V$ can be reduced to inverting $\Gamma$, whose coefficients are bounded in the [1,1][-1,1] interval. We now discuss a metric that can be used to quantify the conditioning of $\Gamma$.
@@ II.1 Correlation dynamic range (CDR)
$\kappa(M)=\frac{\max_{j}(\sigma_{j})}{\min_{j}(\sigma_{j})}\,$, (10)
where the $\sigma_{j}$, $0\leq j<N_{t}$, are the eigenvalues of $M$. The CDR of $M$ is then defined by the value of $\kappa(M)$ in decibels ($\mathrm{d}\mathrm{B}$), namely
@@
The CDR gives the number of orders of magnitude between the smallest and largest eigenvalues of $M$, multiplied by $10$. A general rule of thumb in interpreting the CDR is as follows: a CDR of $x\penalty 10000\ \mathrm{d}\mathrm{B}$ implies that the precision of the data will be degraded by $x/10$ significant digits when multiplied by $M$.
@@ II.2 General data transformations
In this paper, we will investigate a unique class of linear transformation, the regulated Laplace Filters, presented in the next section. However, the general idea in this section may be of use for future research on different transformations.
@@ III Representative sample data
Regarding the M0M-mes-ll, F1M-mes-lh, and F1M-mes-hh correlation functions, we consider standard operators with the quantum numbers of pseudoscalar mesons of the form
@@ IV Improvement of time correlations
In practice, one rarely needs to fit the entire time range of a lattice correlator. Instead, fits generally focus on a limited time range. It is therefore useful to define a restriction operator $R_{S}$, where $S$ is a subset of $\{0,\dots,N_{t}-1\}$, that projects $C$ onto the components indexed by $S$. If $S$ is an interval $\{t_{i},\dots,t_{f}-1\}$, we define $R_{t_{i},t_{f}}=R_{S}$. It is clear that $R_{S}$ is a linear operator. We additionally define the following shorthand notation
@@ IV.3 Summary of correlation improvements
Figure 5: Behavior of the eigenvalues $\sigma_{i}$ of the correlation matrix of the M0M-mes-ll (upper half) and F1M-mes-lh (lower half) correlation functions as a function of Laplace filtering regulator $\lambda$. In each half, the top and bottom rows show the spectrum for the full and restricted time ranges, respectively. In the right column, an additional downsampling step is applied. The faint horizontal lines correspond to the spectrum when no filtering is applied.
In fig.˜5 we illustrate how the individual eigenvalues $\sigma_{i}$ of the correlation matrix vary as a function of $\lambda$ for the M0M-mes-ll and F1M-mes-lh correlation functions. The vertical dashed lines correspond to the values that minimise the CDR for the full (top) or range-restricted correlation matrix. The faint horizontal lines correspond to the unfiltered spectrum. We clearly observe the expected asymptotic behaviour of recovering the original spectrum for very large values of $\lambda$ as well as the emergence of a zero mode when considering the full correlation matrix and very small values of $\lambda$. In table˜2, we summarize the CDRs obtained by filtering and downsampling the M0M-mes-ll and F1M-mes-lh correlators, for both the full time range and a restricted, representative fit range. In all cases, a combination of downsampling and optimized filtering improves the condition number of the correlation matrix by at least four orders of magnitude. This is an encouraging result, and we emphasize that this improvement is achieved solely through a linear, invertible transformation of the data and does not introduce any additional systematic error in subsequent fits of the correlators.
@@ V Spectral effect of Laplace filters
Figure 6: Representation of the pion correlator M0M-mes-$l$ 2-state test $p$-value $p_{\Lambda}$ defined in eq.˜55, with $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$. The four different plots correspond to various values of the initial time $t_{i}$. In all cases, the final time is fixed to $t_{f}=48$. Values below $1\%$ are not represented (2-state hypothesis rejected at $99\%$ confidence level), and values above $1\%$ are within the solid black contour. Values between $1\%$ and $5\%$ are represented in red, and values above $5\%$ are within the dashed contour, and represented by a gradient between orange and green.
Figure 7: Representation of the heavy-light correlator F1M-mes-lh 2-state test $p$-value $p_{\Lambda}$ defined in eq.˜55, with $\Lambda=(\tilde{E}_{0},\tilde{E}_{1})$. The four different plots correspond to various values of the initial time $t_{i}$. In all cases, the final time is fixed to $t_{f}=36$. Plotting conventions are identical to fig.˜6.
@@ V.1.2 Discrete and periodic time
eq.˜42 is the discrete and periodic equivalent of the Källén-Lehmann spectral representation eq.˜34. For $t\in\{0,\dots,N_{t}-1\}$, the expression above can be reduced to the more convenient form
$g_{0}[t;E]=\frac{e^{-Et}+e^{-E(N_{t}-t)}}{1-e^{-N_{t}E}}\,.$ (44)
For $t\neq 0\ (\mathrm{mod}\ N_{t})$, we have
@@
An important conclusion from eq.˜46 is that the Laplace filter can entirely remove the contribution of a given state from a correlator if the regulator is tuned to the energy $\tilde{E}$ associated with this state. On the one hand, this means that when choosing the regulator, one must be careful to avoid the energies of states of interest. On the other hand, it also implies that the filter can be exploited to remove unwanted states, typically excited-state contamination from hadronic interpolating operators.
@@ V.2 Applications to statistical data analysis
Throughout this section, we assume that the correlator under study is given by eq.˜42, with $N_{s}$ states of nonzero amplitude and distinct energies indexed by $n\in\{0,\dots,N_{s}-1\}$. For $t\notin\{-1,0,1\}\ (\mathrm{mod}\ N_{t})$, a single application of the Laplace filter yields eq.˜46. For a vector of $r$ regulators $\Lambda=(\lambda_{0},\dots,\lambda_{r-1})$, we define
@@
i.e., $\smash{C_{\Lambda}}$ is the correlator transformed by $r$ successive Laplace filters. eq.˜46 can be generalized to multiple filters. For $t\notin\{-r,\dots,r\}\ (\mathrm{mod}\ N_{t})$, one obtains
@@
For the remainder of this discussion, we assume that $t\notin\{-r,\dots,r\}\ (\mathrm{mod}\ N_{t})$. In order for the formula above to be valid for at least one time slice, we require $r<N_{t}/2$. Then, eq.˜49 offers an opportunity to determine the energy spectrum $\smash{\tilde{E}_{n}}$ by finding a sequence of Laplace filters that annihilate $C[t]$, i.e., that solve the equation
@@
for all $t$ in a given range satisfying the previous constraints, and containing at least $r$ elements. If the number of filters equals the number of states, i.e., $r=N_{s}$, then eq.˜50 is satisfied if and only if, for all $j\in\{0,\dots,N_{s}-1\}$,
@@
In this case, the energies $E_{n}$ are uniquely determined [^3] This is a consequence of the fact that functions of the form $t\mapsto e^{-Et}$ are linearly independent for distinct values of $E$. by eq.˜50. If $r>N_{s}$, then necessarily the spectrum is a subset of the solution set of eq.˜50, i.e., $\smash{\{\tilde{E}_{0},\dots,\tilde{E}_{N_{s}-1}\}\subset\{\lambda_{0},\dots,\lambda_{r-1}\}}$. If $r<N_{s}$, then eq.˜50 has no solution.
@@
    The number of states $N_{s}$ is given by the minimal $r$ for which eq.˜50 admits a solution;
@@
    The spectrum $\tilde{E}_{j}$ is given by the unique solution of eq.˜50 for $r=N_{s}$.
The procedure above can serve as the basis for new methods to determine the energy spectrum of lattice correlators. Although equation eq.˜50 determines the spectrum uniquely, in practice with noisy simulation data it can only be solved up to statistical noise, introducing ambiguities in the above procedure. Therefore, in what follows, we discuss potential hypothesis-testing methods to constrain the spectrum of the correlator.